OpenStudy (castiel):

sqrt(x+2)<4-x ? solution is -2<=x<2 how?

OpenStudy (jdoe0001):

what did you get? I assume there's no absolute value involved?

OpenStudy (jdoe0001):

anyhow... .so \(\bf \sqrt{x+2} < 4-x\qquad \textit{raise to }{\color{red}{ 2}} \\ \quad \\ (\sqrt{x+2})^2<(4-x)^2\implies x+2 < (4-x)^2 \\ \quad \\ \implies x+2 < (4-x)(4-x)\implies x+2 < 16-4x-4x+x^2 \\ \quad \\ x+2<16-8x+x^2\implies 0 < 14-9x+x^2\implies 0 < x^2-9x+14\) now if you factor the right-side.... see what you get though it's not going to give you \(\bf -2\le x < 2\)

OpenStudy (castiel):

I know, first I did it the same as you wrote but then in solutions it's −2≤x<2. I checked on wolfram and −2≤x<2 is correct. I have no idea how they got that.

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

I'm thinking the "7" answer drops off

OpenStudy (jdoe0001):

lemme do some checking

OpenStudy (anonymous):

domain of \(\sqrt{x+2}\) is \(x\geq 2\)

OpenStudy (anonymous):

ooops i meant \[x\geq -2\]

OpenStudy (anonymous):

in any case whatever you get for the polynomial inequality, you are restrained by \(x\geq -2\) throughout the whole problem

OpenStudy (jdoe0001):

yeap, indeed

OpenStudy (jdoe0001):

@Castiel as you can see, by what satellite73 just pointed out your solutions are correct is 0 < (x-7) (x-2) which makes it 7 < x and 2 < x however, the 7 answers drops off due to DOMAIN RESTRICTIONS on the radical of the original equation "x" cannot be less than -2, for if it does, the radicand turns negative and the radical will give an imaginary value

OpenStudy (castiel):

Thank you

OpenStudy (jdoe0001):

yw